OpenStudy (anonymous):

I have this question: Solve 2|z-2|=36 and came out with {20,-16}. I dont think that is accurate

OpenStudy (ash2326):

2|z-2|=36 or |z-2|=18 if z-2 >=0 then z-2=18 or z=20 if z-2<0 then -(z-2)=18 z=-16 so z=-16 or 20

OpenStudy (cathyangs):

that is correct, Though I don't think that answer in that way is right.

OpenStudy (hoblos):

z-2 =18 =>z=20 OR z-2=-18=> z=-16

OpenStudy (anonymous):

Well my only choices are {20,-16} {20,16} {-20,16} {-20,-16}

OpenStudy (cathyangs):

...normally that's how you would right a set...like domain or range. I guess it's acceptable to give your answer that way, but, like hoblos said, I think 20 or -16 is the standard way.

OpenStudy (anonymous):

Got it thank you.Math homework gets confusing.

OpenStudy (mathteacher1729):

{20, 16} means "the set containing +20 and -16. this is the correct answer. The function y = 2|z-2| intersects the line y = 36 in two places... at x = -16 and at x = +20. (see attached).

OpenStudy (anonymous):

Ohh i see, thank you

OpenStudy (anonymous):

But when i work it out i still get the set containing {20,16} instead of {20,-16}

OpenStudy (anonymous):

Ohh i get it now, i see

OpenStudy (anonymous):